OpenStudy (anonymous):

The fenton college cougars made 40 field goals in a recent basketball game. Some 2 pointers and the rest 3 pointers. all together 40 baskets counted for 89 points. How many of each type of field goals was made.

OpenStudy (anonymous):

31 2 pointers and 9 3 pointers

OpenStudy (anonymous):

How did you get that

OpenStudy (anonymous):

I'm sure there is a way that I just can't remember right now. But I just tried to put in some numbers. I know, not the best and most ideal way :)

OpenStudy (anonymous):

Field goals in a basketball game?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I'll ignore that. ok so this is a systems of equations problem. Your first equation deals with total number of baskets lets call 2 pointers x and 3 pointers y x+y=40 your second equation deals with total points 2x+3y=89 now you have a system of equations to solve x+y=40 2x+3y=89

OpenStudy (anonymous):

the substitution is where im getting hung up on .

OpenStudy (anonymous):

solve for y in the first one y=-x+40 plug that expression into the second one for y and solve for x 2x+3(-x+40)=89

OpenStudy (anonymous):

this is sad to say but im still lost. I don't understand math at all

OpenStudy (anonymous):

you cant solve for x in that equation?

OpenStudy (anonymous):

honestly no. I have been trying to get examples but no luck. I am dumb when it comes to this stuff.

OpenStudy (anonymous):

2x-3x+120=89 -x=-31 x=31

OpenStudy (anonymous):

would it be - 2 +40

OpenStudy (anonymous):

X is 31. since you know x you can find y 31+y=40 y=?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ty